a boy runs 60 yards at a rate of 10miles an hour how long did it take him to finish

Answer by stanbon(57967)

a boy runs 60 yards at a rate of 10miles an hour how long did it take him to finish
---------
Time = distance/rate = (60/1760 )miles)/(10 mph) = (6/1760) hrs
-------
Cheers,
Stan H.